OpenStudy (anonymous):

PLEASE HELP!...I know how to solve this problem..but my issue is if I have to convert 7.3 C to Kelvin! A sample of silver absorbed 9.86 J of heat, and its temperature increased from 23.2°C to 30.5°C. What was the mass of the silver sample? (The specific heat of silver is 0.235 Jg-1K-1.)

OpenStudy (rogue):

To convert from Celsius to kelvin, use this\[K = 273.15 + C\]After you do that, to solve your problem, use \[Q = mc \Delta T\]Good luck!

OpenStudy (rogue):

The celsius to kelvin conversion really isn't needed for this problem, since the difference between the temperatures (delta T) will still be the same, but you can do it for practice =)

OpenStudy (anonymous):

i know how to convert from Celsius to kelvin .....so you are saying that I don't need to convert it? that is my question.

OpenStudy (anonymous):

So I can keep 7.3 in Celsius

OpenStudy (anonymous):

?

OpenStudy (kainui):

Nope, 7.3 is in Celsius AND Kelvin. You might be wondering how this is possible. Well, you see 7.3 is the CHANGE in temperature, not a temperature itself. To see what I mean, calculate the change before and after you convert the temperatures to Kelvin. 30.5 C -23.2 C = 7.5 C 303.65 K - 296.35 = 7.5 K It doesn't matter if you're using Celsius or Kelvin since the change in temperature is exactly the same amount. It doesn't work for Fahrenheit because a degree of F is different, since 1 degree of C is the same as 1 degree of K. I hope that helps!

OpenStudy (anonymous):

THANK YOU!! THAT MAKES SENSE :D!!....Also is that a flag of Israel in your profile pic?

OpenStudy (kainui):

Nope, it's a sailboat, I love sailing. Haha, glad I could help.